# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v2

@@ V-E Reinforcement Learning with Human Feedback
As machines and their associated systems become increasingly complex, the conditions that need to be considered maybe take a long time for Algorithm 5. Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data. By integrating human feedback into the RL process, the framework can:
@@
Thus, we further incorporate the following information into our proposed framework in Algorithm 5:
@@
    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states from the C-MAPSS dataset.
@@
    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriateness of its actions.
@@
based on which, the modified reward function for DRL is
@@
1 Initialize the neural network weights;
@@ 16       Periodically adjust the learning process based on accumulated human feedback;
Algorithm 5 The RLHF Algorithm for Maintenance Actions Selection
V-F Results for RLHF
@@
Figure 15: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
@@
Figure 16: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
@@
Figure 15 and Figure 16 reflect the changes after adding on the human feedback in the reward function. The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized. This is affected by the weight of human feedback in the reward function as well as the complexity of the task. Since our task converges very fast even without human feedback, there is not much room for improvement. In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
VI Conclusions
@@
In this paper, we present a comprehensive prescriptive maintenance framework that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations. The framework starts by employing the Transformer network to estimate the RUL distribution of engines, offering a robust and adaptive data-driven approach to understanding machine conditions in real-time. As machine operates, our model continues to dynamically updating RUL estimates, providing an ever-evolving status of the equipment’s condition.
We rigorously compare the performance of several state-of-the-art time-series forecasting models, including TFT, NBeats, and DeepAR, to validate the accuracy of our RUL estimates. Our findings indicate that the Transformer-based algorithm delivers superior predictive performance, laying a solid foundation for the subsequent decision-making process. For DRL-based maintenance action recommendation, we demonstrate the efficiency of SAC, DQN, and PPO. Each of these algorithms has its strengths and weaknesses, but they all are adapted to working seamlessly with the dynamically updated RUL estimates. Future work may involve integrating more complex cost models, incorporating uncertainty estimates, and testing the framework in a multi-agent setting or across different types of machines.
In conclusion, our work signifies a step forward in the request for more intelligent, adaptive, and cost-effective prescriptive maintenance systems. By leveraging cutting-edge algorithms for both RUL estimation and action recommendation, we offer a solution that dynamically adapts to the operational realities, thereby enhancing both economic and operational efficiencies.

### v3

@@ IV-D Integrate with Human Feedback for Maintenance Action Recommendation
Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods. Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively. Thus, as machines and their associated systems become increasingly complex, conditions that need to be considered may take a long time for Algorithm 4. Human experts, with their working experiences and domain knowledge to find complex patterns, can provide valuable insights that may not be readily available through raw data. To minimize labor costs, we assume that one expert will be involved during the human feedback process. By integrating human feedback into the RL process, the framework can:
@@
The above goals are achieved by incorporating the following information into our proposed framework in Algorithm 5:
@@
    Human Opinions: Collect human predictions of the desired behavior. Experienced engineers can provide opinions on maintenance decisions based on various machine states obtained from the C-MAPSS dataset.
@@
    Feedback on Agent’s Actions: After the agent makes a decision, a human can provide feedback on the chosen action, guiding the agent on the appropriate actions.
@@
based on which, the modified reward function from (IV-C3) for DRL becomes
@@
1
2Initialize the neural network weights;
@@ 22       Periodically adjust the learning process based on accumulated human feedback;
Algorithm 5 A RLHF-based algorithm for optimal maintenance action recommendation
V Case Study
In this section, we present a case study of our proposed TranDRL framework.
V-A Data
NASA’s C-MAPSS dataset [26] contains four subsets denoted as FD001, FD002, FD003, and FD004, each with a training and test set, reflecting distinct operating and fault modes. The training data comprise inputs from $3$ operating conditions and $21$ sensor measurements across four engines of the same type. While the datasets are i.i.d, each engine has its unique operational settings and sensor readings. By leveraging the C-MAPSS dataset, we can gain insights into how IoT systems might perform in real-world scenarios, particularly those involving complex machinery such as aircraft engines. The dataset facilitates data collection and monitoring, mimicking real-world engine performance, which represents the general working scenario for IoT-based sensing and decision making in different industrial sectors. Therefore, it provides a convenient platform for the development and testing of predictive maintenance and anomaly detection algorithms by offering simulated engine degradation and fault data.
V-B Experiment Setup
In the experiment, we applied a piecewise linear degradation model [27] for RUL determinations. Each engine’s last cycle data point generally stands as the sample, with actual RUL values included in the dataset.
Environment Settings. We utilized an Intel Core i9-14900K CPU processor, 128GB DDR5 RAM, and a PNY RTX 4090 GPU card. The system ran on Ubuntu 22.04 LTS with Python 3.9 and PyTorch 2.0.1 for tasks.
Hyperparamters. We carefully choose its corresponding hyperparameters based on empirical experiments.
Transformers. The TFT model employs a learning rate of 0.03, hidden layers of size 16, a single attention head, a dropout rate of 0.1, a hidden continuous size of 8, and Symmetric Mean Absolute Percentage Error (SMAPE) as the loss function, focusing on the target variable RUL. We also introduce two baseline algorithms for prediction performance comparison. For the first reference algorithm, the N-Beats model [28], also set with a learning rate of 0.03, uses an encoder length of 120, a prediction length of 60, a batch size of 128, implement early stopping at 10 epochs, and uses gradient clipping at 0.1. For another reference algorithm, the DeepAR model [29], parameters include a lower learning rate of 0.003, encoder and decoder lengths of 120 and 60 respectively, the same batch size of 128, target normalization via GroupNormalizer, training up to 10 epochs with early stopping after 4, and weight decay at 0.01. The negative binomial distribution loss is employed.
Deep Q-network (DQN). DQN reduces the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function [30]. It also introduces experience replay to improve learning stability and efficiency. The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences. We use a batch size of $256$ and update the target network every $50$ step. The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we use a neural network with two hidden layers of $256$ units each, and we utilize the Adam optimizer with a learning rate of $10^{-4}$.
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy [31]. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size of $256$. Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $10^{-4}$. Finally, the GAE parameter is set to $0.95$ to estimate the advantage function.
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces [32]. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks is set to $10^{-4}$, and we use a discount factor of $0.99$. Then, we apply a temperature parameter $\alpha$ of $0.2$ to balance the entropy term in the objective function. The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$. We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$. Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $10^{-4}$.
(a) The RUL prediction results for machine FD001.
(b) The RUL prediction results for machine FD002.
(c) The RUL prediction results for machine FD003.
(d) The RUL prediction results for machine FD004.
Figure 2: RUL prediction results for four machines.
Baseline. The ground truth is obtained from run-to-failure data, where the machine is operated until failure occurs. The run-to-failure data consists of time-series measurements of various sensor data recorded during operation until failure. The ground truth RUL values are calculated by subtracting the current cycle $t$ from the cycle $T_{f}$ at which failure occurred, expressed mathematically as $\text{RUL}(t)=T_{f}-t$. For baseline methods, we utilize NBeats [28] and DeepAR [29] to compare against our proposed method, aiming to verify whether it outperforms the baselines.
V-C Results and Analysis
Figures 2(a)-2(d) demonstrate the superior performance of our proposed algorithm (Algorithm 2 as TFT) over NBeats and DeepAR in predicting the RUL for an individual machine, as evidenced by consistently higher RUL values. This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in high dimensional sensor data, highlighting its robustness across various machine behaviors.
(a) Comparison of RUL prediction results of four machines using the model trained from machine FD001.
(b) Comparison of RUL prediction results of four machines using the model trained from machine FD002.
(c) Comparison of RUL prediction results of four machines using the model trained from machine FD003.
(d) Comparison of RUL prediction results of four machines using the model trained from machine FD004.
(e) Comparison of RUL prediction results of four machines using the global model.
Figure 3: Comparison of RUL prediction results of four machines using the model.
In contrast, Figure 3(e) shows that while the federated learning models (Algorithm 3) generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models. However, Figures 3(a) to 3(d) show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments. This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
(a) The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
(b) The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 4: Results of Using Deep Reinforcement Learning for Optimizing Maintenance Strategies.
@@ V-C Results and Analysis
(a) The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
@@
(b) The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms with human feedback in our designed environment.
Figure 5: Results for Using Reinforcement Learning with Human Feedback for Optimizing Maintenance Strategies.
We compare the performance of DQN, PPO, and SAC in the framework of TranDRL (Algorithm 4) in our designed maintenance task in Figures 4(a) and 4(b). The evaluation metric is the total reward obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving a higher total reward on the task. The average total reward obtained by PPO is $55,000$, while DQN and SAC achieve an average total reward of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC. In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed.
However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster. Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve optimal performance.
@@
Figure 5(a) and Figure 5(b) reflect the changes after adding on the human feedback in the reward function (Algorithm 5). The results show that the speed of convergence does not improve too much, but the process becomes more stable and variances minimized. This is affected by the weight of human feedback in the reward function as well as the complexity of the task. Since our task converges very fast even without human feedback, there is not much room for improvement. In the future, we will conduct experiments on more complex real-world scenarios to confirm the effectiveness of the RLHF algorithm.
VI Conclusion and Future Work
@@
In this paper, we present a prescriptive maintenance framework (TranDRL) that integrates real-time RUL estimations with DRL algorithms for dynamic action recommendations. The framework starts by employing the Transformer network to estimate the RUL distribution of machines (e.g., engines), offering a robust and adaptive data-driven approach to understanding machine conditions in real-time. As the machine operates, our model continues to dynamically update RUL estimates, providing an ever-evolving status of the equipment’s condition. We rigorously compare the performance of several state-of-the-art time-series forecasting models, including NBeats and DeepAR, with our proposed method to validate the accuracy of our RUL estimates. For the framework of DRL-based maintenance action recommendations, we demonstrate the efficiency of SAC, DQN, and PPO with human feedback. Each of these algorithms has its strengths and weaknesses, but they are all adapted to working seamlessly with the dynamically updated RUL estimates.
The limitations of our framework include its inability to process image-based data and the lack of metrics to evaluate the accuracy of human experts’ contributions. These shortcomings may compromise the performance of the framework. Future work will focus on refining the framework, such as introducing new neural network designs, to address these limitations.